OpenStudy (anonymous):

0.5n + 17 + n = 20 (Show Work)

OpenStudy (lgbasallote):

add all the n's and transpose the non n's 0.5n + n = 20 - 17 1.5n = 3 n = 2

OpenStudy (anonymous):

can you show work please @Igbasallote